Experimental HEP Computing

Experimental HEP group runs workstation “cluster”

• SGI Crimson (4D-CRIM) file server (R4000) – 8 Gb disk

– tape drive

– FDDI and Thinwire ethernet • 3 Indigo II workstations (R4400)

– 11 and 14 Gb of SCSI disk – Exabyte tape drives

– FDDI and Thinwire

• SGI 4D-35 workstation (R3000)

• Two processors of Irides 4D-280 (R3000)

• Six older DEC stations

• Seven X-terminals

• Sun SPARC with 2 Gb disk

• 5 PC’s for secretary and lab use

Computing Resource Needs

Group has relied on PCS to provide:

• System software maintenance (Varsity)

• System hardware maintenance and diagnosis – FDDI loop

– Peripherals

– Monitors

• Backup of critical data – Want to redefine set of disks

• E-mail and network access

• Consultation/assistance on hardware purchases – expedite orders

– install new peripherals

• access to modem pool and printers

• Consultation on security issues

New Projects

Have a number of outstanding projects:

• Installation and testing of MBONE – Do we have necessary bandwidth to FNAL/CERN?

• Become ASF client site for HEP software

• Complete unification of cluster – unified password file

– common login directories from all machines

– sharing of all disks and tapes?